26,28,31,35,?

26,28,31,35,?

Grade:11

2 Answers

Sher Mohammad IIT Delhi
10 years ago
40


sher mohammad
B.Tech, IIT Delhi

Anant Jain
10 years ago
Ans is 40 because 26+2=28 28+3=31 31+4=35 35+5=40 this means in every number positive integer started from 2 is added to number to get next number